OpenStudy (anonymous):

Let . If F is an antiderivative of the same function f and , find I'll post the question again

OpenStudy (anonymous):

\[Let \int\limits_{0}^{5}f(x) dx= 19 . If F is an antiderivative of the same function f and F(0)=13 , find F(5) .\]

OpenStudy (anonymous):

If F is an antiderivative of the same function f and F(0)=13 , find F(5)

OpenStudy (anonymous):

An integral is the antiderivative evaluated at the endpoints, in this case 0 and 5, so it could be rewritten as F(5)-F(0)=19. So F(5)-13=19, and F(5)=32.

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

thanks a ton

OpenStudy (anonymous):

No problem, glad it helped

OpenStudy (anonymous):

One more question